# CustomVideo: Customizing Text-to-Video Generation with Multiple Subjects

### v1

@@ 4.1 Experimental Setup
We collect a dataset CustomStudio with 69 individual subjects and construct 57 meaningful pairs for multi-subject driven T2V generation. These subjects are adapted from DreamBooth Ruiz et al. (2023), CustomDiffusion Kumari et al. (2023) and internet. The objects cover a wide range of categories, including pet, person, toy, transportation, scene, plant, wearable item, and so on. Each pair of subjects has 10 different text prompts, which are designed with different contexts, actions, and so on.
@@
We train CustomVideo for 500 steps with AdamW Loshchilov and Hutter (2017) optimizer (batch size 2, learning rate 4e-5, and weight decay 1e-2). For class-specific prior preservation, we collect 200 class images from LAION-400M Schuhmann et al. (2021) for each subject. Note that the corresponding class images are also concatenated during training phase. The negative value $\eta$ in the mask $\boldsymbol{\mathcal{M}}$ is set as -1e-8. The weight parameters $\alpha$ and $\beta$ in Eq. (4) are set as 0.1 and 1.0, respectively. During inference, we perform 50 steps denoising with DDIM Song et al. (2020) sampler and classifier-free guidance Ho and Salimans (2022). The resolution of generated 24-frame videos are $576\times 320$ with 8 fps. Interestingly, we find that the trained weights and word tokens learned from low resolution zeroscope T2V model can be directly loaded for high resolution one [^2] https://huggingface.co/cerspense/zeroscope_v2_XL to generate personalized videos with $1024\times 576$ resolution, in which no additional training computation cost is required. Our CustomVideo is implemented based on Diffusers von Platen et al. (2022). The training phase takes about 8 minutes for a subject pair on 1 RTX 3090 GPU. Meanwhile, it takes about 1 minute and 3 minutes to generate a low and high resolution video on 1 RTX 3090 GPU, respectively.
@@
Except for VideoDreamer Chen et al. (2023b), we also consider adapting previous SOTA image-based multi-subject driven methods to a video scenario for comparison, including DreamBooth Ruiz et al. (2023) and CustomDiffusion Kumari et al. (2023). DreamBooth trains all of the weights in a diffusion model to personalize the generation from multiple subjects. CustomDiffusion fine-tune the diffusion model with joint training on multiple subjects. VideoDreamer proposes a disen-mix and human-in-the-loop fine-tuning strategy for multi-subject customization of T2V model. For fair comparison, we implement all the comparison method and our CustomVideo based on zeroscope model.
@@
Table 1: Quantitative results of our CustomVideo with comparison to SOTA methods, including DreamBooth, CustomDiffusion, and VideoDreamer. Our proposed CustomVideo consistently outperforms previous SOTA methods for all 4 evaluation metrics.
@@
DreamBooth 0.6476 0.6015 0.3098 0.7209
CustomDiffusion 0.6581 0.6218 0.3186 0.7548
VideoDreamer 0.6649 0.6314 0.3471 0.7362
CustomVideo (ours) 0.7051 0.6749 0.3955 0.8142
Evaluation Metrics.
Following previous works Wei et al. (2023); Chen et al. (2023b), we quantitatively evaluate our CustomVideo with the following 4 metrics: 1) CLIP Textual Alignment computes the average cosine similarity between the generated frames and text prompt with CLIP Radford et al. (2021) ViT-B/32 Dosovitskiy et al. (2021) image and text models; 2) CLIP Image Alignment calculates the average cosine similarity between the generated frames and subject images with CLIP ViT-B/32 image model; 3) DINO Image Alignment measures the average visual similarity between generated frames and reference images with DINO Caron et al. (2021) ViT-S/16 model; 4) Temporal Consistency Esser et al. (2023) evaluates the average cosine similarity of between all consecutive frame pairs in the generated videos.
4.2 Main Results
Quantitative Results.
We conduct quantitative experiments on our collected dataset. To ensure a thorough analysis, we generate videos using 10 individual prompts and 4 random seeds for each pair of subjects, resulting in a total of 2,280 generated videos for each method. We then evaluate the quality of the generated videos using four metrics, and the results are presented in Table 1. The table clearly demonstrates that our proposed method is capable of generating videos that are better aligned with the given subjects. This improvement can be attributed to our specially designed co-occurrence and attention control mechanisms, which effectively disentangle and preserve the fidelity of the subjects. Notably, our CustomVideo outperforms the most recent method, VideoDreamer, by 6.05%, 6.89%, and 13.94% in terms of CLIP Textual Alignment, CLIP Image Alignment, and DINO Image Alignment, respectively. Furthermore, our CustomVideo generates videos with significantly higher temporal consistency compared to state-of-the-art methods, as indicated in Table 1. For instance, CustomVideo surpasses VideoDreamer by 10.59% in terms of Temporal Consistency.
Figure 4: User study. Our CustomVideo achieves the best human preference compared with 3 SOTA comparison methods in terms of Textual Alignment, Image Alignment, and Overall Quality.
Table 2: Quantitative results for component analysis of our proposed CustomVideo. We observe a significant performance drop when removing subjects concatenation or background removal. In practice, we suggest simultaneously using positive and negative attention mechanisms to obtain the best results.
w/o remove bg 0.6676 0.6248 0.3498 0.7891
w/o concat 0.6189 0.6371 0.3572 0.7936
both single and concat 0.6148 0.6319 0.3208 0.7983
w/o pos. attn. 0.6858 0.6402 0.3569 0.8015
w/o neg. attn. 0.6926 0.6392 0.3694 0.8075
CustomVideo (ours) 0.7051 0.6749 0.3955 0.8142
Qualitative Results.
We present the qualitative comparison results in Figure 3. From these results, we observe that our CustomVideo approach effectively ensures the co-occurrence of multiple subjects and successfully disentangles different subjects. However, both DreamBooth and CustomDiffusion fail to capture the structural color information of the car provided in our experiments. The dominant black color from the car window obscures the entire car, resulting in low fidelity. Moreover, the generated frames from VideoDreamer lack consistency, as some frames depict two cars while others only show one car. Additionally, VideoDreamer fails to capture accurate color information, such as the color of the car door. In contrast, our CustomVideo method excels in handling such challenging scenes and foreground subject scenarios, effectively capturing the intricate structural details of the provided car for video generation. Similarly, when considering the case of ‘cat’ and ‘dog’, our approach also demonstrates superior capability in generating high-quality videos.
Human Preference Study.
To further validate our method, we conduct human evaluations on our CustomVideo with comparison to 3 SOTA methods. In this study, we collect 900 answers from 25 independent human raters with the following questions: 1) which one is aligned to the text prompt best? 2) which one is aligned to the subject images best? 3) which one has the best overall quality? The results are shown in Figure 4. It can be found that our proposed CustomVideo is the most preferred one from the human side from all these three dimensional evaluation.
4.3 Ablation Studies
We conduct ablation analysis of our method from 2 aspects: 1) the effect of each component in our method; 2) the effect of hyper-parameters in attention control mechanism.
Table 3: Quantitative results of the weight of attention loss. In practice, $\alpha$=0.1 yields the best result and is adopted in all experiments.
@@ 4.3 Ablation Studies
1.0 0.6784 0.6651 0.3792 0.8015
0.1 0.7051 0.6749 0.3955 0.8142
0.01 0.6892 0.6584 0.3817 0.8092
Table 4: Quantitative results of the negative value. Interestingly, $\eta$=-1e-8 achieves the best result in terms of all the considered metrics. Practically, we adopt this setting in all the experiments.
@@
-1e-5 0.6891 0.6582 0.3917 0.7996
-1e-8 0.7051 0.6749 0.3955 0.8142
-1e-11 0.6918 0.6627 0.3895 0.8079
Component Analysis.
We conduct a thorough analysis of each component in our CustomVideo, presenting both qualitative and quantitative results in Figure 5 and Table 2. These findings provide valuable insights into the importance and effectiveness of each component. One crucial observation is the significance of ensuring co-occurrence during the fine-tuning process. We notice a significant drop in performance when multiple subjects are not concatenated into a single image (lines ‘w/o concat’ and ‘CustomVideo’ in Table 2). The absence of subject concatenation leads to the domination of one single subject during fine-tuning, which is deemed unacceptable (line ‘w/o concat’ in Figure 5). Moreover, we conduct an ablation study on fine-tuning the diffusion model with both single subjects and concatenated subjects to examine if the single subject could aid in learning corresponding characteristics. Surprisingly, the results reveal that adding a single subject to the training process proved detrimental, resulting in inconsistent generation during inference (line ‘both single and concat’ in Figure 5). Furthermore, we observe that removing the background from the given subject images significantly improves the generation of high-quality videos. By eliminating the background, the diffusion model can focus solely on learning the characteristics of the given subjects. The comparison between videos generated with and without background removal clearly illustrates the monotonous nature of videos lacking this component (line ‘w/o remove bg’ in Figure 5). Our attention control mechanism plays a crucial role in aligning the generated videos with the given subjects. The positive attention guidance in particular improves the CLIP Image Alignment metric by 5.42%, as demonstrated in Table 2. The negative attention guidance also proved beneficial in promoting better image alignment. Moreover, the attention control mechanism significantly enhances the temporal consistency of the generated videos, a crucial factor in video generation. The comparison between videos generated without positive attention guidance and videos generated without negative attention guidance clearly showcases the positive impact of both components (lines ‘w/o pos. attn.’ and ‘w/o neg. attn.’ in Figure 5). The positive guidance specifically preserves the unique characteristics of a subject, such as distinguishing between a cat and a dog, while the negative guidance weakens the influence of other subjects on a specific subject. By incorporating both positive and negative guidance, our CustomVideo excels in generating videos with high subject fidelity and remarkable temporal consistency.
Figure 6: Failure cases of our CustomVideo. Our CustomVideo may fail to capture some nuanced details of subjects when meeting extremely similar pairs of subjects, such as ‘sloth’ and ‘teddybear’ (first line). Also, our approach fails to generate vivid facial contents with a global view, such as ‘man’ and ‘bike’ (second line)
Effect of Hyper-parameters in Attention Control.
In addition to component analysis, we also conduct investigations into the effects of two important parameters: the weight of attention loss ($\alpha$) and the negative value ($\eta$) used in the guidance mask ($\boldsymbol{\mathcal{M}}$). The quantitative results of these investigations are presented in Table 3 and Table 4. Regarding the weight of attention loss, we observe that the best performance is achieved when $\alpha$ is set to 0.1. This value results in optimal alignment between the generated videos and the given subjects, indicating the importance of appropriately balancing the attention guidance during training. For the negative value in the guidance mask, our findings reveal that even a slight negative value is sufficient to enhance the quality of T2V generation. This implies that incorporating negative attention guidance can effectively suppress the influence of other subjects on a specific subject, improving the generation quality.
@@ 4.4 Limitations
In Figure 6, we present some failure cases encountered during our experiments. One challenge arises when generating subjects that are extremely similar, such as a sloth and a teddy bear, where capturing fine-grained details becomes difficult. Additionally, since our method relies on the ability of base model, it would fail if the base model could not generate, such as small faces. To address the failure cases, we can improve CustomVideo by using a stronger T2V base model, like Stable Video Diffusion Blattmann et al. (2023a). Additionally, we can implement mechanisms to prioritize and emphasize nuanced characteristics in the given images during training.
@@ 5 Conclusions
This paper provides a novel framework CustomVideo for multi-subject driven T2V generation, powered by a simple yet effective co-occurrence and attention control mechanism. To evaluate the proposed framework, we collect a comprehensive dataset consisting of 69 individual subjects and 57 meaningful pairs. Extensive experiments and human preference study on collected dataset demonstrate the effectiveness of CustomVideo, outperforming existing methods significantly.
Appendix A Additional Ablation Studies
Layers of Cross Attention Maps.
We investigates the effect of how we extract the cross attention maps from the cross attention layers. There are four levels of cross attention layers in the UNet, in which the sizes cross attention maps are different. Taking the resolution $576\times 320$ of our generated low resolution video as an example, the sizes of cross attention maps are $72\times 40$ ($\mathcal{\ell}_{1}$), $36\times 20$ ($\mathcal{\ell}_{2}$), $18\times 10$ ($\mathcal{\ell}_{3}$), and $9\times 5$ ($\mathcal{\ell}_{4}$). Here, we study the effect of different levels of cross attention maps, the results are shown in Table 5. We observe a significant performance drop with too large ($\mathcal{\ell}_{1}$) or small ($\mathcal{\ell}_{4}$) cross attention maps. Regarding the cross attention maps with middle size from $\mathcal{\ell}_{2}$ and $\mathcal{\ell}_{3}$, $\mathcal{\ell}_{3}$ works better. Thus, in our experiments, we extract cross attention maps from $\mathcal{\ell}_{3}$ for attention control.
Appendix B CustomStudio Dataset

### v2

@@ 4.1 Experimental Setup
We collect a dataset CustomStudio with 63 individual subjects and construct 68 meaningful pairs for multi-subject driven T2V generation. These subjects are adapted from DreamBooth [45], CustomDiffusion [40], and Mix-of-Show [52], in which they cover a wide range of 13 categories. Each pair of subjects has 10 different text prompts for evaluation, which are designed with different contexts, actions, and so on. More details of dataset can be found in Sec. A.
@@
We train CustomVideo for 500 steps with AdamW [53] optimizer (batch size 2, learning rate 4e-5, and weight decay 1e-2). For class-specific prior preservation, we generate 200 class images with Stable Diffusion v2.1 [13] for each subject, in which the generating prompts are obtained from Claude-3-Opus [54]. Note that the corresponding class images are also concatenated during training phase. The negative value $\eta$ in the mask $\boldsymbol{\mathcal{M}}$ is set as -1e-8. The weight parameters $\alpha$ and $\beta$ in Eq. (4) are set as 0.2 and 1.0, respectively. During inference, we perform 50 steps denoising with DPM-Solver [8] sampler and classifier-free guidance [55]. The resolution of generated 24-frame videos are $576\times 320$ with 8 fps. Interestingly, we find that the trained weights and word tokens learned from low resolution T2V model can be directly loaded for high resolution zeroscope T2V model [56] to generate personalized videos with $1024\times 576$ resolution, in which no additional training computation cost is required. Our CustomVideo is implemented based on Diffusers [57]. The training phase takes about 8 minutes for a subject pair on 1 RTX 3090 GPU. Meanwhile, it takes about 1 minute and 2 minutes to generate a low and high resolution video on 1 RTX 3090 GPU, respectively.
@@
Except for most relevant VideoDreamer [12], we also consider adapting previous SOTA image-based multi-subject driven methods to a video scenario for comparison, including DreamBooth [45] and CustomDiffusion [40]. Also, we compare our method with single-subject driven video personalization method DreamVideo [11]. Moreover, we consider an alternative way for multi-subject T2V customization, that is, the combination of multi-concept T2I and I2V diffusion models. For this case, we utilize the latest multi-concept driven T2I method DisenDiff [14] for customization and I2V method SVD-XT [15] for animating the generated image.
Figure 3: Qualitative results of our CustomVideo with comparison to SOTA methods, including DreamBooth, CustomDiffusion, VideoDreamer, DisenDiff + SVD-XT, and DreamVideo. The first line indicates the given subjects, while each line indicates the frames generated by the corresponding method. The bottom line is the text prompt for inference. We can observe that our CustomVideo can generate videos with better fidelity of subjects compared with previous SOTA methods.
Evaluation Metrics.
Following previous works [11, 12], we quantitatively evaluate our CustomVideo with the following 4 metrics: 1) CLIP Textual Alignment computes the average cosine similarity between the generated frames and text prompt with CLIP [58] ViT-B/32 [59] image and text models; 2) CLIP Image Alignment calculates the average cosine similarity between the generated frames and subject images with CLIP ViT-B/32 image model; 3) DINO Image Alignment measures the average visual similarity between generated frames and reference images with DINO [60] ViT-S/16 model; 4) Temporal Consistency [61] evaluates the average cosine similarity of all consecutive frame pairs in the generated videos.
4.2 Main Results
Qualitative Results.
We present the qualitative comparison results in Figure 3. From these results, we observe that our CustomVideo approach effectively ensures the co-occurrence of multiple subjects and successfully disentangles different subjects. However, both DreamBooth and CustomDiffusion fail to capture the structural color information of the car provided in our experiments. The dominant black color from the car window obscures the entire car, resulting in low fidelity. Moreover, the generated frames from VideoDreamer lack consistency, as some frames depict two cars while others only show one car. Additionally, VideoDreamer fails to capture accurate color information, such as the color of the car door. The alternative approach, DisenDiff + SVD-XT can not neither retain the identities of the given subjects nor generate videos with rich motions. Without specific design for tackling multiple subjects, DreamVideo can not distinguish the similar pair ‘cat’ and ‘dog’. In contrast, our CustomVideo method excels in handling such challenging scenes and foreground subject scenarios, effectively capturing the intricate structural details of the provided car for video generation. Similarly, when considering the case of ‘cat’ and ‘dog’, our approach also demonstrates superior capability in generating high-quality videos. More results can be found in Figure 9 and Figure 10.
Method Reference CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ T. Cons. $\uparrow$
DreamBooth [45] CVPR’23 0.6451 0.6079 0.3109 0.7084
CustomDiffusion [40] CVPR’23 0.6524 0.6206 0.3164 0.7342
VideoDreamer [12] arXiv’23 0.6638 0.6297 0.3479 0.7267
DisenDiff [14] + SVD [15] CVPR’24 0.6592 0.6179 0.3318 0.7095
DreamVideo [11] CVPR’24 0.6646 0.6128 0.3228 0.7498
CustomVideo (ours) / 0.7075 0.6863 0.3983 0.7960
@@ 4.2 Main Results
Table 1: Quantitative results of our CustomVideo with comparison to SOTA methods, including DreamBooth, CustomDiffusion, VideoDreamer, DisenDiff + SVD-XT, and DreamVideo. Our proposed CustomVideo consistently outperforms previous SOTA methods for all 4 evaluation metrics.
Quantitative Results.
We conduct quantitative experiments on our collected dataset. To ensure a thorough analysis, we generate videos using 10 individual prompts and 4 random seeds for each pair of subjects, resulting in a total of 2,720 generated videos for each method. We then evaluate the quality of the generated videos using four metrics, and the results are presented in Table 1. The table clearly demonstrates that our proposed method is capable of generating videos that are better aligned with the given subjects, outperforming the most recent video diffusion model, DreamVideo, by 11.99% and 23.39% in terms of CLIP Image Alignment and DINO Image Alignment, respectively. Meanwhile, the videos generated from our approach achieve better alignment with the desired text prompts, exceeding DreamVideo by 6.46% CLIP Textual Alignment. These improvements can be attributed to our specially designed co-occurrence and attention control mechanisms, which effectively disentangle and preserve the fidelity of the subjects. Furthermore, our CustomVideo generates videos with significantly higher temporal consistency compared to SOTA methods, as indicated in Table 1. For instance, CustomVideo surpasses DreamVideo by 6.16% in terms of Temporal Consistency.
Figure 4: User study. Our CustomVideo achieves the best human preference compared with 5 SOTA comparison methods in terms of Textual Alignment, Image Alignment, and Overall Quality.
Human Preference Study.
To further validate our method, we conduct human evaluations on our CustomVideo with comparison to 5 SOTA methods. In this study, we collect 1500 answers from 25 independent human raters with the following questions: 1) which one is aligned to the text prompt best? 2) which one is aligned to the subject images best? 3) which one has the best overall quality? The results are shown in Figure 4. Our CustomVideo is found to be the most preferred option based on human evaluations in all three dimensions.
4.3 Ablation Studies
We conduct ablation analysis of our method from 3 aspects: 1) the effect of each component in our method; 2) the behavior of attention control; 3) the effect of hyper-parameters in attention control.
Component Analysis.
We conduct a thorough analysis of each component in our CustomVideo, presenting both qualitative and quantitative results in Figure 5 and Table 3. One crucial observation is the significance of ensuring co-occurrence during the fine-tuning process. We notice a significant drop in performance when multiple subjects are not concatenated into a single image (line ‘w/o concat’ in Table 3). The absence of subject concatenation leads to the domination of one single subject, which is deemed unacceptable (line ‘w/o concat’ in Figure 5). Moreover, we conduct a study on fine-tuning T2V model with both single and concatenated subjects to examine if the single subject could aid in learning corresponding characteristics. Surprisingly, the results reveal that adding the single subjects to the training process proved detrimental, resulting in inconsistent generation during inference (line ‘both single and concat’ in Figure 5).
Our attention control mechanism plays a crucial role in preserving the identities of subjects. The positive attention guidance in particular improves the CLIP Image Alignment metric by 11.45%, as demonstrated in Table 3. The negative attention guidance also proved beneficial in promoting better image alignment. Moreover, the attention control mechanism significantly enhances the temporal consistency of the generated videos, a crucial factor in video generation. The videos generated without positive and negative attention guidance clearly showcase the positive impact of both components (lines ‘w/o pos. attn.’ and ‘w/o neg. attn.’ in Figure 5). The positive guidance specifically preserves the unique characteristics of a subject, such as distinguishing between a cat and a dog, while the negative guidance weakens the influence of other subjects on a specific subject. By incorporating both positive and negative guidance, our CustomVideo excels in generating videos with high subject fidelity and remarkable temporal consistency. Furthermore, we observe that removing the background from the given subject images significantly improves the generation of high-quality videos (line ‘w/o remove bg’ in Figure 5). By eliminating the background, the T2V model can focus solely on learning the characteristics of the given subjects.
Comparison of Cross-attention Maps.
We investigate the cross-attention maps during training, and study the behavior of our attention control mechanism. The visualization results are shown in Figure 6. Without attention control, it can be observed that the learnable word token ‘<new1>’ for cat is aligned to the region of dog, which is unacceptable. In contrast, equipped with attention control mechanism, the learnable word token ‘<new1>’ and ‘<new2>’ are better aligned to the correct areas of cat and dog, respectively. Such results strongly demonstrate the effectiveness of our approach.
Figure 6: Comparison results of cross-attention maps. We observe that the learnable word tokens better match the corresponding correct areas of the subjects with our proposed attention control.
Effect of Hyper-parameters in Attention Control.
We also conduct investigations into the effects of two important parameters: the weight of attention loss ($\alpha$) and the negative value ($\eta$) used in the guidance mask ($\boldsymbol{\mathcal{M}}$). The quantitative results of these investigations are presented in Table 3a and Table 3b. We observe that the best performance is achieved when the weight $\alpha$ is set to 0.2. This value results in optimal alignment between the generated videos and the given subjects, indicating the importance of appropriately balancing the attention guidance during training. For the negative value in the guidance mask, our findings reveal that even a slight negative value is sufficient to enhance the quality of T2V generation. This implies that incorporating negative attention guidance can effectively suppress the influence of other subjects on a specific subject, improving the generation quality.
@@ 4.3 Ablation Studies
w/o remove bg 0.6682 0.6752 0.3636 0.7609
w/o concat 0.6115 0.6611 0.3386 0.7034
both single and concat 0.6248 0.6096 0.3054 0.7604
w/o pos. attn. 0.6851 0.6158 0.3215 0.7213
w/o neg. attn. 0.6961 0.6716 0.3613 0.7886
CustomVideo (ours) 0.7075 0.6863 0.3983 0.7960
Table 2: Quantitative results for component analysis of our proposed CustomVideo. We observe a significant performance drop when removing subjects concatenation or background removal. Moreover, training the model with both single and concatenated subjects will produce unstable results, leading to performance degradation. In practice, we suggest simultaneously using positive and negative attention mechanisms to obtain the best results.
@@
1.0 0.6909 0.6496 0.3773 0.7812
0.2 0.7075 0.6863 0.3983 0.7960
0.01 0.6924 0.6710 0.3715 0.7886
(a) Weight $\alpha$ for attention loss
@@
-1e-5 0.6786 0.6392 0.3719 0.7604
-1e-8 0.7075 0.6863 0.3983 0.7960
-1e-11 0.6926 0.6737 0.3613 0.7812
(b) Negative value $\eta$ for GT mask
Table 3: Quantitative results of the weight of attention loss and the negative value. Interestingly, $\alpha$=0.2 and $\eta$=-1e-8 yield the best results. Practically, we adopt this setting in all the experiments.
Figure 7: Failure cases of our CustomVideo. Our CustomVideo fails to generate vivid facial contents with a global view, such as ‘man’ and ‘bike’ (first line). Also, our approach can not tackle too many subjects, e.g., customizing T2V generation with 6 different subjects (second line).
@@ 4.4 Limitations
In Figure 7, we present some failure cases encountered during our experiments. Since our method relies on the ability of the base model, it would fail if the base model could not generate, such as small faces in a global view. Currently, Our approach can not tackle too many subjects, in which the spatial positions of different objects in the video conflict with each other. To address such case, we can control the positions of different subjects via conditioning on the pre-defined spatial bounding boxes during inference sampling.
@@ 5 Conclusions
This paper provides a novel framework CustomVideo for multi-subject driven T2V generation, powered by a simple yet effective co-occurrence and attention control mechanism. It successfully disentangles different similar subjects and preserves the corresponding identities. We design a flexible and efficient training process, only requiring subject images rather than subject videos. And during inference, a desired video can be easily generated just by providing a text prompt from the user. We collect a comprehensive dataset CustomStudio for evaluation. Extensive quantitative and qualitative experiments, together with a user study, demonstrates the effectiveness of our CustomVideo, outperforming SOTA methods significantly. We hope our method could serve as a strong baseline and motivate further research on subject-driven applications, especially for multi-subject scenarios.
Figure 8: The overview of our CustomStudio dataset. Samples in our CustomStudio dataset cover a wide range of 13 object categories, including anime character, decorate item, furniture, instrument, person, pet, plant, plush, scene, thing, toy, transport, and wearable item.
Appendix A CustomStudio Dataset
Dataset Info.